# Empowering Biomedical Discovery with AI Agents

### v1

@@
We envision “AI scientists” as systems capable of skeptical learning and reasoning that empower biomedical research through collaborative agents that integrate machine learning tools with experimental platforms. Rather than taking humans out of the discovery process, biomedical AI agents combine human creativity and expertise with AI’s ability to analyze large datasets, navigate hypothesis spaces, and execute repetitive tasks. AI agents are proficient in a variety of tasks, including self-assessment and planning of discovery workflows. These agents use large language models and generative models to feature structured memory for continual learning and use machine learning tools to incorporate scientific knowledge, biological principles, and theories. AI agents can impact areas ranging from hybrid cell simulation, programmable control of phenotypes, and the design of cellular circuits to the development of new therapies.
1 Introduction
A long-standing ambition for artificial intelligence (AI) in biomedicine is the development of AI systems that could eventually make major scientific discoveries, with the potential to be worthy of a Nobel Prize—fulfilling the Nobel Turing Challenge [1]. While the concept of an “AI scientist” is aspirational, advances in agent-based AI pave the way to the development of AI agents as conversable systems capable of skeptical learning and reasoning that coordinate large language models (LLMs), machine learning (ML) tools, experimental platforms, or even combinations of them [2, 3, 4, 5] (Figure 1).
@@ 1 Introduction
The complexity of biological problems requires a multistage approach, where decomposing complex questions into simpler tasks is necessary. AI agents can break down a problem into manageable subtasks, which can then be addressed by agents with specialized functions for targeted problem-solving and integration of scientific knowledge, paving the way toward a future in which a major biomedical discovery is made solely by AI [2, 6]. In the near future, AI agents can accelerate discovery workflows by making them faster and more resource-efficient. Looking ahead, AI agents can enable insights that might not have been possible using ML alone by making predictions across temporal and spatial scales prior to experimental measurements at those scales, and can eventually identify new modes of behavior within biological systems [6].
@@
This vision is possible thanks to advances in LLMs [7, 8, 9], multimodal learning, and generative models. Chat-optimized LLMs, such as GPT-4 [10], can incorporate feedback, enabling AI agents to cooperate through conversations with each other and with humans [11]. These conversations can involve agents providing and seeking advice, critique, and validation [12, 13]. Then, since a single LLM can exhibit a broad range of capabilities—especially when configured with appropriate prompts and inference settings, conversations between differently configured agents can combine these capabilities in a modular manner [14]. LLMs have also demonstrated the ability to solve complex tasks by breaking them into subtasks [15, 16]. However, suppose we follow conventional approaches to foundation models such as LLMs and other large pre-trained models. In that case, we may not develop AI agents that can generate novel hypotheses because such novelty would not have been in the data used to train the model, suggesting that current foundation models alone are not sufficient for “AI scientists”. Using LLMs as a comparison, generating novel hypotheses requires creativity and grounding in scientific knowledge, whereas generating novel text requires adherence to semantic and syntactic rules [17], so the latter approach aligns well with techniques for next-token prediction within LLMs, while the former does not.
@@
Here, we offer a perspective that “AI scientists” can be realized as AI agents backed by humans, LLMs, ML models, and other tools like experimental platforms that cooperate to solve complex tasks. An AI agent should be skeptical when developing biological hypotheses, capable of characterizing its uncertainty and using that as a driver to acquire and refine its scientific knowledge bases in a way that human scientist partners can trust [18]. AI agents should be designed to adapt to new biological insights, incorporate the latest scientific findings, and refine hypotheses based on experimental results. This adaptability ensures agents remain relevant in the face of rapidly evolving biological data and understanding [17]. However, the challenge lies in developing algorithms that allow for this continual learning without suffering from catastrophic forgetting [19], which requires a balance between acquiring new knowledge and retaining existing information. By addressing this challenge, AI agents can power platforms, such as closed-loop synthesis frameworks, autonomous microscopy facilities, and hybrid cell simulators.
@@
Ethical considerations arise from using AI agents [20, 21]. Allowing AI agents to make changes in environments through tools or calls to experimental platforms can be dangerous. Safeguards and protocols need to be in place to prevent harm and negative outcomes [22]. Conversely, discovery workflows might include conversations between AI agents (but no interaction with environments is allowed). In that case, we need to consider the impact of such interactions on scientists and their reliance on AI agents. The impact of AI agents on biology, coupled with the challenges they pose, underscores the need for responsible implementation. AI agents have the potential to perform research and operational tasks under the oversight of humans [2].
2 Evolving use of data-driven models in biomedical research
@@
Data-driven models have reshaped biomedical research over the past several decades through the development of databases, search engines, machine learning, and interactive and foundation learning models (Figure 2). These models have advanced modeling of proteins [23, 24, 25, 26, 27], genes [28], phenotypes [29], clinical outcomes [30, 31, 32], and chemical compounds [33, 34] through mining of biomedical data.
@@
Databases and search engines. In biological research, databases (DBs)  [35, 36, 37] aggregate knowledge from experiments and studies, offering searchable repositories containing standardized biological data vocabularies. An example of such a database is the AlphaFold Protein Structure DB [38], which includes more than 200 million protein structures predicted by AlphaFold [39]. Molecular search engines retrieve information from these databases [40, 41, 42]. FoldSeek [43] retrieves protein structures from the AlphaFold DB by translating query structures into 3D interaction alphabet sequences and using pretrained substitution matrices.
@@
Distinct from search engines, designed for retrieving information based on specific queries, AI agents are capable of reasoning to formulate search queries and subsequently acquire information. Curated databases offer structured and factual information, aiding in reducing the risks associated with misinformation potentially generated by agent hallucinations [44]. A notable feature of these agents is their ability to dynamically retrieve information when needed and to create and reflect upon the obtained passages. This reflection process renders the agent controllable during inference, allowing for customization of its actions to meet the requirements of tasks beyond what is possible using search engines and database queries.
@@
Machine learning models. Beyond information retrieval, ML models excel in identifying patterns and assimilating latent knowledge to generalize predictions about novel data [45, 46, 47]. An example is AlphaFold [39] protein folding model that uses multi-sequence alignment with a deep learning model to predict the 3D protein structure from amino acid sequences, achieving near-atom level accuracy. AI agents represent an evolution in ML models, building on the foundations of successes such as the transformer architecture [48] and generative pretraining [9]. These agents have reasoning and interactive capabilities that distinguish them from ML models, which typically require specialized models for each task. Unlike traditional ML models, agents assess the evolving environment, which is valuable for modeling dynamic biological systems.
@@
Interactive learning models. Interactive learning, often referred to as active learning [49] and reinforcement learning [50], represents a further advancement in ML models. This method enhances the adaptability and efficiency of ML techniques by incorporating exploration and feedback mechanisms. Active learning focuses on efficiently training models with limited labeled data, proving particularly advantageous in biological learning where data may be insufficient. It selectively queries the most informative data points for labeling and optimizing the learning process, which improves how models learn with data. Reinforcement learning involves an agent learning how to act by observing the results of past actions in an environment, mirroring the trial-and-error approach. In biological research, interactive learning has been applied to various fields, such as small molecule design [51], protein design [52, 53], drug discovery [54, 55], perturbation experiment design [56], and cancer screening [57]. For instance, GENTRL [51] uses reinforcement learning to navigate the chemical space and find candidate chemical compounds that can act against a biological target.
Leveraging interactive learning, AI agents achieve greater autonomy in information retrieval tasks. Active learning improves training efficiency through data labeling selected to maximize model performance. However, AI agents extend beyond this data-centric approach; for example, reinforcement learning with human feedback (RLHF) [50] uses a “reward model” to train and fine-tune an LLM-based agent with direct human feedback to understand human instruction naturally.
@@
AI agents. Looking ahead, AI agents have advanced capabilities, including proactive information acquisition through perception modules, interaction with tools, reasoning, and the capability to engage with and learn from their environments. Agents use external tools, such as lab equipment, and have perception modules, such as integrated visual ML tools to receive information from the environment (Section 5). Agents can incorporate search engines and ML tools and process information across data modalities via perception modules to generate hypotheses and refine them based on scientific evidence [2, 3].
3 Types of biomedical AI agents
@@
The prevailing approach to building agents nowadays is to use LLMs, where a single LLM is programmed to perform various roles (Section 3.1). However, beyond LLM agents, we envision multi-agent systems for discovery workflows that combine heterogeneous agents (Figure 1) consisting of ML tools, domain-specific specialized tools, and human experts (Section 3.2). Given that much of biomedical research is not text-based, such agents have broader applicability to biomedicine.
3.1 Large language model based AI agents
@@
Programming a single LLM with diverse roles equips LLM-based agents with conversational interfaces that emulate human expertise and can access tools [58, 59] (Figure 3a). The rationale behind this approach stems from pretraining an LLM to encode general knowledge, followed by in-domain fine-tuning of the LLM to encode domain-specific specialist knowledge and aligning the LLM with human users through role-playing and conversation. Instruction tuning [60] can be used for the former by training the LLM to follow human instruction through the use of prompt examples, including dialogues that incorporate biological reasoning [61]. Additionally, RLHF optimizes LLM performance by selecting the most human-preferred outputs from a range of responses to specific prompts, further aligning LLMs with human roles. Consequently, a single LLM, programmed to fulfill multiple roles, can provide a more practical and effective solution than developing specialized models. By assigning specific roles, the agents can replicate the specialized knowledge of experts across various fields, such as structural biology, genetics, and chemistry, surpassing the capabilities of querying a non-specialized LLM [62] and performing tasks previously not possible [63]. Early results in clinical medicine question-answering suggest that assigning specific roles to GPT-4 [62] can achieve better performance than using domain-specialized LLMs like BioGPT [64], NYUTron [65], and Med-PaLM [66, 67].
@@
We envision three approaches for assigning roles to biological AI agents: domain-specific fine-tuning, in-context learning, and automatic generation of optimized roles. The first approach involves instruction-tuning an LLM across a large number of biological tasks to ground the LLM in the biological domain, followed by RLHF to ensure that the tuned LLM performs tasks aligned with the goals, wants, and needs of scientists. The second approach uses in-context learning of LLMs [68] to process longer contextual information provided in inputs, such as biologist-generated instructions, enabling agents to grasp the domain context for each task. This approach is supported by using textual prompts to define agent roles [63, 69]. Both strategies require biologists to carefully gather task-specific data or craft prompts. However, as roles defined by humans may not always direct agents as intended, there has been a movement towards allowing LLM-based agents more autonomy in role specification. This paradigm shift in role definition enables agents to autonomously generate and refine role prompts, engaging in self-directed learning and role identification. For instance, Fernando et al. [70] demonstrate an agent’s ability to evolve and tailor its prompts in reaction to user inputs. Similarly, Yang et al. [71] investigate the application of LLM as an optimizer to enhance prompt refinement and optimization for improved performance in assigned roles. Through this self-referential learning framework, agents transition from task executors to entities capable of more autonomous learning.
@@
The agent system, comprising a single LLM prompted to adopt various roles, has proven to be an effective auxiliary in scientific research. Studies suggest that agents allocated specific roles exhibit enhanced capabilities compared to either sequentially querying a single LLM or employing a single tool repetitively. A case in point is Coscientist [2], which shows the potential of GPT-4-based agents for chemical research tasks, including optimizing reactions for palladium-catalyzed cross-couplings. Within Coscientist, GPT-4 undertakes the role of a planner, serving as a research assistant. The agent uses in-context prompts to support the use of tools such as web and documentation search, code execution via Python API, and even symbolic lab language (SLL) [2]. To complete tasks that require access to a physical device, the planning agent starts with a prompt provided by the scientist and uses search tools to compile the documentation for the experiment. Following this, the agent generates SLL code and executes it, which entails transferring it onto the device and controlling the device.
3.2 Multi-agent AI systems
@@
Brainstorming agents (Figure 3b). Brainstorming research ideas with multiple agents constitutes a collaborative session to generate a broad spectrum of research concepts through the joint expertise of scientists and agents. In such sessions, agents are prompted to contribute ideas, prioritizing the volume of contributions over their initial quality to foster creativity and innovation. This method encourages the proposal of unconventional and novel ideas, allowing participants to build upon the suggestions of others to uncover new avenues of inquiry while withholding judgment or critique. The process enables agents to apply their domain knowledge and resources to form a collective idea pool. This pool can then be distilled and examined more thoroughly. To foster divergent thinking and creativity, it can be beneficial for agents to specialize in areas of biology, for instance, microglia biology, neuronal degeneration, and neuroinflammation in the case of a multi-agent system for Alzheimer’s.
Expert consultation agents (Figure 3c). Expert consultation entails soliciting expertise from individuals or entities with specialized knowledge. This process involves expert agents gathering information from various sources and providing insights, solutions, decisions, or evaluations in response. Other agents or humans then refine their approaches based on this feedback. Liang et al. [73] illustrate that LLMs can function like human expert reviewers, offering scientific critiques on research manuscripts. Similarly, an AI agent might consult another agent specialized in a specific area to refine ideas within AI systems, mirroring the mentor-mentee dynamics found in academic environments. In another example, in addressing Alzheimer’s and related dementias, diagnosing Alzheimer’s based on cognitive criteria might present borderline cases. Consulting an AI agent could offer additional perspectives, determining if such cases align with Alzheimer’s based on brain pathology or alternative biomarkers.
@@
Round table discussion agents (Figure 3e). Round-table discussions involve multiple agents engaging in a process that fosters the expression of diverse viewpoints to make collaborative decisions on the topics under discussion. In such sessions, agents articulate their ideas and insights, pose questions, and provide feedback on others’ contributions. They then respond to these queries, refine their initial propositions based on feedback, or attempt to persuade their peers. This method promotes equal participation among all agents, urging them to contribute their expertise and perspectives, offer constructive criticism, question underlying assumptions, and suggest amendments to improve the proposed solutions. For instance, Reconcile [74] implements a collaborative reasoning scheme among LLM agents through successive rounds of dialogue. Agents attempt to convince each other to adjust their responses and use a confidence-weighted voting mechanism to achieve a more accurate consensus than if a single LLM-based agent is used. During each discussion round, Reconcile orchestrates the interaction between agents using a ‘discussion prompt,’ which includes grouped answers and explanations produced by each agent in the preceding round, their confidence levels, and examples of human explanations for correcting answers.
@@
Self-driving lab agents (Figure 3f). The self-driving laboratory is a multi-agent system where the end-to-end discovery workflow is iteratively optimized under the broad direction of scientists but without requiring step-by-step human oversight [75]. Once the agent system is trained, it can describe experiments necessary to test the generated hypotheses, analyze the results of said experiments, and use them to improve its internal scientific knowledge models. Agents in the self-driving system need to address the following three elements: determine inductive biases to reduce the search space of hypotheses, implement methods to rank order hypotheses considering their potential biomedical value with experimental cost, characterize skepticism via uncertainty quantification and analysis of experiments in reference to the original hypothesis, and refine hypotheses using data and counterexamples from experiments [76]. Ideally, hypothesis agents are creative and skeptical when developing biological hypotheses that extrapolate indirectly from the existing body of knowledge [17]. Experimental agents steer operational agents that use a combination of in silico approaches and physical platforms to execute experiments. Following this, reasoning agents integrate the latest results to guide future experimental design. The utility of experimental results, such as the yield of high-throughput screening of a chemical library against a biological target, can be compared for different versions of the agent system given a time budget for hypothesis and experiment generation.
4 Levels of autonomy in AI agents
@@
AI agents, when integrated with experimental platforms, can operate at varying levels of autonomy tailored to the diverse requirements across biomedical fields. We classify these AI agents into four levels according to their proficiency in three areas of discovery: Hypothesis, Experiment, and Reasoning (Table 1). Specific capabilities within each area define these levels, necessitating that agents exhibit the capabilities for a given level across all areas (an agent with Level 3 capabilities in the Experiment area but only Level 2 capabilities in Reasoning and Hypothesis areas would be classified as Level 2).

### v2

@@
We envision “AI scientists” as systems capable of skeptical learning and reasoning that empower biomedical research through collaborative agents that integrate AI models and biomedical tools with experimental platforms. Rather than taking humans out of the discovery process, biomedical AI agents combine human creativity and expertise with AI’s ability to analyze large datasets, navigate hypothesis spaces, and execute repetitive tasks. AI agents are poised to be proficient in various tasks, planning discovery workflows and performing self-assessment to identify and mitigate gaps in their knowledge. These agents use large language models and generative models to feature structured memory for continual learning and use machine learning tools to incorporate scientific knowledge, biological principles, and theories. AI agents can impact areas ranging from virtual cell simulation, programmable control of phenotypes, and the design of cellular circuits to developing new therapies.
Introduction
@@
A long-standing ambition for artificial intelligence (AI) is the development of AI systems that can eventually make major scientific discoveries, learn on their own, and acquire knowledge autonomously. While this concept of an “AI scientist” is aspirational, advances in agent-based AI pave the way to the development of AI agents as conversable systems capable of reflective learning and reasoning that coordinate large language models (LLMs), machine learning (ML) tools, experimental platforms, or even combinations of them [1, 2, 3, 4] (Figure 1). The complexity of biological problems requires an approach where a complex problem is decomposed into simpler tasks. AI agents can break down a problem into manageable subtasks, which can then be addressed by agents with specialized functions for targeted problem-solving and integration of scientific knowledge [1, 5]. In the near future, AI agents can accelerate discovery workflows by making them faster and more resource-efficient. AI agents improve the efficiency of routine tasks, automate repetitive processes, and analyze large datasets to navigate hypothesis spaces at a scale and precision that surpasses current human-driven efforts. This automation allows for continuous, high-throughput research that would be impossible for human researchers to perform alone at the same scale or speed. Looking further ahead, AI agents can enable insights that might not have been possible using ML alone by making predictions across temporal and spatial scales prior to experimental measurements at those scales and can eventually identify new modes of behavior within biological systems [5].
@@
This vision is possible thanks to advances in LLMs [6, 7, 8], multimodal learning, and generative models. Chat-optimized LLMs, such as GPT-4 [9], can incorporate feedback, enabling AI agents to cooperate through conversations with each other and with humans [10]. These conversations can involve agents seeking human feedback and critique, and identifying gaps in their knowledge [11, 12]. Then, since a single LLM can exhibit a broad range of capabilities—especially when configured with appropriate prompts and inference settings, conversations between differently configured agents can combine these capabilities in a modular manner [13]. LLMs have also demonstrated the ability to solve complex tasks by breaking them into subtasks [14, 15]. However, suppose we follow conventional approaches to foundation models such as LLMs and other large pre-trained models. In that case, we may not develop AI agents that can generate novel hypotheses because such novelty would not have been in the data used to train the model, suggesting that current foundation models alone are not sufficient for “AI scientists”. Using LLMs as a comparison, generating novel hypotheses requires creativity and grounding in scientific knowledge, whereas generating novel text requires adherence to semantic and syntactic rules [16], so the latter approach aligns well with techniques for next-token prediction within LLMs, while the former does not.
@@
Here, we offer a perspective that “AI scientists” can be realized as AI agents backed by humans, LLMs, ML models, and other tools like experimental platforms that form a compound AI system. An AI agent should be able to formulate biomedical hypotheses, critically evaluate them, characterize their uncertainty, and use that as a driver to acquire and refine its scientific knowledge bases in a way that human scientists can trust [17]. AI agents should be designed to adapt to new biological insights, incorporate the latest scientific findings, and refine hypotheses based on experimental results. This adaptability ensures agents remain relevant in the face of rapidly evolving biological data [16], balancing between encoding new findings and retaining old knowledge [18].
Realizing this perspective shift, biomedical AI agents can impact areas ranging from virtual cell simulation, programmable control of phenotypes, and the design of cellular circuits to developing new therapies. Virtual cell simulation involves creating detailed models of cellular processes, where AI can predict the effects of genetic modifications or drug treatments on cell behavior. This can allow for an understanding of cellular mechanisms and generation of testable hypotheses, reducing the time and cost of traditional methods. Programmable control of phenotypes leverages AI agents to design precise genetic modifications to study gene functions. For example, CRISPR-based gene editing guided by an AI agent can activate or inhibit specific genes across large cell populations in a multi-round editing campaign. Each round involves identifying the next edit based on the user-specified target phenotype and experimental readout from the previous round. Designing cellular circuits involves using AI agents to predict the behavior of genetic components and optimize their arrangement to create circuits that perform tasks such as sensing environmental changes or producing therapeutic proteins.
@@
Ethical considerations arise from biomedical AI agents [19, 20]. Allowing them to make changes in environments through ML tools or calls to experimental platforms can be dangerous. Safeguards need to be in place to prevent harm [21]. Conversely, discovery workflows might include conversations between AI agents (but no interaction with environments is allowed). In that case, we need to consider the impact of such interactions on scientists and their reliance on AI agents. Further, challenges uniquely relevant to biomedical AI agents include the lack of large experimental datasets that cover diverse use cases beyond the current focus on a handful of biomedical domains like structural biology and single-cell science. AI agents need to represent biomedical knowledge in a data-efficient manner and achieve strong generalization to new tasks with little or no additional training. Biomedical AI agents can assist with research and operations under human oversight, but their impact and challenges highlight the need for responsible implementation.
Evolving use of data-driven models in biomedical research
@@
Over the past several decades, data-driven models have reshaped biomedical research by developing databases, search engines, machine learning, and interactive and foundation learning models (Figure 2). These models have advanced modeling of proteins [22, 23, 24, 25, 26], genes [27], phenotypes [28], clinical outcomes [29, 30, 31], and chemical compounds [32, 33] through mining of biomedical data.
@@
Databases and search engines. In biological research, databases (DBs)  [34, 35, 36] aggregate knowledge from experiments and studies, offering searchable repositories containing standardized biological data vocabularies. An example of such a database is the AlphaFold Protein Structure DB [37], which includes more than 200 million protein structures predicted by AlphaFold [38]. Molecular search engines retrieve information from these databases [39, 40, 41]. FoldSeek [42] retrieves protein structures from the AlphaFold DB by translating query structures into 3D interaction alphabet sequences and using pretrained substitution matrices. Search engines are designed to retrieve information based on specific queries, lacking the ability to refine these queries through reasoning. They cannot iteratively process obtained information to refine results or customize subsequent actions. Additionally, while databases reduce the risk of misinformation through curated data, they lack mechanisms to identify and remove irrelevant information.
@@
Distinct from search engines, AI agents are capable of reasoning to formulate search queries and subsequently acquire information. Curated databases offer structured and factual information, aiding in reducing the risks associated with misinformation potentially generated by agent hallucinations [43, 44]. For example, the retrieval-augmented generation [44] is equipped for AI agents to answer questions based on scientific literature. A notable feature of these agents is their ability to retrieve information when needed and to create and iteratively process the obtained passages. This reflection process makes the agent controllable during inference, allowing for customization of its actions to meet task requirements beyond what is possible using search engines and database queries.
@@
Machine learning models. Beyond information retrieval, ML models excel in identifying patterns and assimilating latent knowledge to generalize predictions about novel data [45, 46]. Existing machine learning models typically require specialized models for each task and do not possess the reasoning and interactive capabilities that distinguish AI agents. An example is the AlphaFold [38], which predicts 3D protein structures with high accuracy using multi-sequence alignment with a deep learning model but is tailored for protein folding. AI agents represent an evolution in ML models, building on the foundations of successes such as the transformer architecture [47] and generative pretraining [8]. These agents’ reasoning and interactive capabilities distinguish them from ML models, which typically require specialized models for each task. Unlike traditional ML models, agents assess the evolving environment, which is valuable for modeling dynamic biological systems.
@@
Interactive learning models. Interactive learning, often referred to as active learning [48] and reinforcement learning [49], represents a further advancement in ML models by incorporating exploration mechanisms and human feedback. Active learning strategies can help build models for datasets with small sample sizes when conventional ML models might be insufficient due to limited statistical power. It selectively queries the most informative data points for labeling and optimizing the learning process, which improves how models learn with data. Reinforcement learning involves an agent learning how to act by observing the results of past actions in an environment, mirroring the trial-and-error approach. In biological research, interactive learning has been used for small molecule design [50], protein design [51, 52], drug discovery [53, 54], perturbation experiment design [55], and cancer screening [56]. For instance, GENTRL [50] uses reinforcement learning to navigate the chemical space and identify chemical compounds that can act against biological targets. However, interactive models are predominantly designed for narrow use cases and struggle to generalize to new goals without retraining the models from scratch. Leveraging interactive learning, AI agents achieve greater autonomy in information retrieval tasks. Active learning improves training efficiency through data labeling selected to maximize model performance. However, AI agents extend beyond this data-centric approach; for example, reinforcement learning with human feedback (RLHF) [49] uses a “reward model” to train and fine-tune an LLM-based agent with direct human feedback to understand human instruction naturally.
@@
AI agents. Biomedical AI agents have advanced capabilities, including proactive information acquisition through perception modules, interaction with tools, reasoning, and engaging with and learning from their environments. Agents use external tools, such as lab equipment, and have perception modules, such as integrated visual ML tools, to receive information from the environment. Agents can incorporate search engines and ML tools and process information across data modalities via perception modules to generate hypotheses and refine them based on scientific evidence [1, 2].
Types of biomedical AI agents
@@
The prevailing approach to building agents is to use LLMs, where a single LLM is programmed to perform various roles. However, beyond LLM agents, we envision multi-agent systems for discovery workflows that combine heterogeneous agents (Figure 1) consisting of ML tools, domain-specific specialized tools, and human experts. Given that much of biomedical research is not text-based, such agents have broader applicability to biomedicine than LLM-based agents alone.
Large language model based AI agents
@@
Programming a single LLM with diverse roles equips LLM-based agents with conversational interfaces that emulate human expertise and can access tools [57, 58] (Figure 3a). The rationale behind this approach stems from pretraining an LLM to encode general knowledge, followed by in-domain fine-tuning of the LLM to encode domain-specific specialist knowledge and aligning the LLM with human users through role-playing and conversation. Instruction tuning [59] can be used for the former by training the LLM to follow human instruction through prompt examples, including dialogues that incorporate biological reasoning [60]. Additionally, RLHF optimizes LLM performance by selecting the most human-preferred outputs from a range of responses to specific prompts, further aligning LLMs with human roles. Consequently, a single LLM, programmed to fulfill multiple roles, can provide a more practical and effective solution than developing specialized models. By assigning specific roles, the agents can replicate the specialized knowledge of experts across various fields, such as structural biology, genetics, and chemistry, surpassing the capabilities of querying a non-specialized LLM [61] and performing tasks previously not possible [62]. Early results in clinical medicine question-answering suggest that assigning specific roles, such as clinicians, to GPT-4 [61] can achieve better performance in terms of accuracy on multiple-choice benchmarks compared to using domain-specialized LLMs like BioGPT [63], NYUTron [64], and Med-PaLM [65, 66].
@@
We envision three approaches for assigning roles to biological AI agents: domain-specific fine-tuning, in-context learning, and automatic generation of optimized roles. The first approach involves instruction-tuning an LLM across many biological tasks to ground the LLM in the biological domain, followed by RLHF to ensure that the tuned LLM performs tasks aligned with scientists’ goals, wants, and needs. The second approach uses in-context learning of LLMs [67] to process longer contextual information provided in inputs, such as biologist-generated instructions, enabling agents to grasp the domain context for each task. This approach is supported by using textual prompts to define agent roles [62, 68]. Both strategies require biologists to carefully gather task-specific data or craft prompts. However, as roles defined by humans may not always direct agents as intended, there has been a movement towards allowing LLM-based agents more autonomy in role specification. This paradigm shift in role definition enables agents to autonomously generate and refine role prompts, engaging in self-directed learning and role identification. For instance, An agent’s ability to evolve and tailor its prompts in reaction to user inputs has been demonstrated in [69]. Additionally, the application of LLM as an optimizer to enhance prompt refinement and optimization for improved performance in assigned roles has been investigated in [70]. Through this self-referential learning framework, agents transition from task executors to entities capable of more autonomous learning.
@@
The agent system, comprising a single LLM prompted to adopt various roles, has shown to be a valuable support tool in scientific research. Studies suggest that agents allocated specific roles exhibit enhanced capabilities compared to either sequentially querying a single LLM or employing a single tool repetitively. A case in point is Coscientist [1], which shows the potential of GPT-4-based agents for chemical research tasks, including optimizing reactions for palladium-catalyzed cross-couplings. Within Coscientist, GPT-4 undertakes the role of a planner, serving as a research assistant. The agent uses in-context prompts to use tools such as web and documentation search, code execution via Python API, and even symbolic lab language (SLL) [1]. To complete tasks that require access to a physical device, the planning agent starts with a prompt provided by the scientist and uses search tools to compile the documentation for the experiment. Following this, the agent generates SLL code and executes it, which entails transferring it onto the device and controlling the device.
Multi-agent AI systems
@@
Brainstorming agents (Figure 3b). Brainstorming research ideas with multiple agents constitutes a collaborative session to generate a broad spectrum of research concepts through the joint expertise of scientists and agents. In such sessions, agents are prompted to contribute ideas, prioritizing the volume of contributions over their initial quality to foster creativity and innovation. This method encourages the proposal of unconventional and novel ideas, allowing participants to build upon the suggestions of others to uncover new avenues of inquiry while withholding judgment or critique. The process enables agents to apply their domain knowledge and resources to form a collective idea pool. Each agent would provide insights and generate hypotheses based on their specialized knowledge, which the group can then integrate and refine. For example, in a multi-agent system designed for Alzheimer’s research, agents could specialize in microglia biology, neuronal degeneration, and neuroinflammation. To explore new therapeutic targets for Alzheimer’s disease, an agent specialized in microglia biology might propose investigating the role of microglial cells in synaptic pruning, while another agent focused on neuronal degeneration could suggest examining the protective effects of certain neurotrophic factors. These diverse ideas are pooled together, allowing researchers to explore a comprehensive range of potential research directions.
Expert consultation agents (Figure 3c). Expert consultation entails soliciting expertise from individuals or entities with specialized knowledge. This process involves expert agents gathering information from various sources and providing insights, solutions, decisions, or evaluations in response. Other agents or humans then refine their approaches based on this feedback. LLMs have the potential to assist in offering scientific critiques on research manuscripts, as demonstrated in recent studies [72]. However, LLMs lack the nuanced understanding of human reviewers and should be seen as complementary to, not a replacement for, human expertise. Similarly, an AI agent might consult another agent specialized in a specific area to refine ideas within AI systems, mirroring the mentor-mentee dynamics found in academic environments. In another example, in addressing Alzheimer’s and related dementias, diagnosing Alzheimer’s based on cognitive criteria might present borderline cases. Consulting an AI agent could offer additional perspectives, determining if such cases align with Alzheimer’s based on brain pathology or alternative biomarkers.
@@
Round table discussion agents (Figure 3e). Round table discussions involve multiple agents engaging in a process that fosters the expression of diverse viewpoints to make collaborative decisions on the topics under discussion. In such sessions, agents articulate their ideas and insights, pose questions, and provide feedback on others’ contributions. They then respond to these queries, refine their initial propositions based on feedback, or attempt to persuade their peers. This method promotes equal participation among all agents, urging them to contribute their expertise and perspectives, offer constructive criticism, question underlying assumptions, and suggest amendments to improve the proposed solutions. For instance, Reconcile [73] implements a collaborative reasoning scheme among LLM agents through successive rounds of dialogue. Agents attempt to convince each other to adjust their responses and use a confidence-weighted voting mechanism to achieve a more accurate consensus than if a single LLM-based agent is used. During each discussion round, Reconcile orchestrates the interaction between agents using a ‘discussion prompt,’ which includes grouped answers and explanations produced by each agent in the preceding round, their confidence levels, and examples of human explanations for correcting answers.
@@
Self-driving lab agents (Figure 3f). The self-driving laboratory is a multi-agent system where the end-to-end discovery workflow is iteratively optimized under the broad direction of scientists but without requiring step-by-step human oversight [74]. Once the agent system is trained, it can describe experiments necessary to test the generated hypotheses, analyze the results of said experiments, and use them to improve its internal scientific knowledge models. Agents in the self-driving system need to address the following three elements: determine inductive biases to reduce the search space of hypotheses, implement methods to rank order hypotheses considering their potential biomedical value with experimental cost, characterize skepticism via uncertainty quantification and analysis of experiments in reference to the original hypothesis, and refine hypotheses using data and counterexamples from experiments [75]. Ideally, hypothesis agents are creative and reflective when developing biological hypotheses that extrapolate indirectly from the existing body of knowledge [16]. There is emerging evidence that generative models have the potential to generate novel hypotheses. [76] demonstrated that using latent knowledge from published materials science literature can recommend novel materials. [77] leveraged LLMs trained with an autoregressive pretraining objective to predict molecules. Experimental agents steer operational agents that use a combination of in silico approaches and physical platforms to execute experiments. Reasoning agents integrate the latest results to guide future experimental design. The utility of experimental results, such as the yield of high-throughput screening of a chemical library against a biological target, can be compared for different versions of the agent system given a time budget for hypothesis and experiment generation.
Levels of autonomy in AI agents
@@
When integrated with experimental platforms, AI agents can operate at varying levels of autonomy tailored to the diverse requirements across biomedical fields. We classify these AI agents into four levels according to their proficiency in three areas of discovery: Hypothesis, Experiment, and Reasoning (Table 1). Specific capabilities within each area define these levels, necessitating that agents exhibit the capabilities for a given level across all areas (an agent with Level 3 capabilities in the Experiment area but only Level 2 capabilities in Reasoning and Hypothesis areas would be classified as Level 2).